89 students are in a class, 7 out of 8 are girls, how many are girls are there?
Answer by stanbon(75887)

89 students are in a class, 7 out of 8 are girls, how many are girls are there?
-----
# of girls = (7/8)*89 = 77 7/8 = 78 when rounded up
=====================
Cheers,
Stan H.
=============